# Landing with the Score: Riemannian Optimization through Denoising

### v1

@@ Abstract
Under the data manifold hypothesis, high-dimensional data are concentrated near a low-dimensional manifold. We study the problem of Riemannian optimization over such manifolds when they are given only implicitly through the data distribution, and the standard manifold operations required by classical algorithms are unavailable. This formulation captures a broad class of data-driven design problems that are central to modern generative AI. Our key idea is to introduce a link function that connects the data distribution to the geometric operations needed for optimization. We show that this function enables the recovery of essential manifold operations, such as retraction and Riemannian gradient computation. Moreover, we establish a direct connection between our construction and the score function in diffusion models of the data distribution. This connection allows us to leverage well-studied parameterizations, efficient training procedures, and even pretrained score networks from the diffusion model literature to perform optimization. Building on this foundation, we propose two efficient inference-time algorithms—Denoising Landing Flow (DLF) and Denoising Riemannian Gradient Descent (DRGD)—and provide theoretical guarantees for both feasibility (approximate manifold adherence) and optimality (small Riemannian gradient norm). Finally, we demonstrate the effectiveness of our approach on finite-horizon reference tracking tasks in data-driven control, highlighting its potential for practical generative and design applications.
@@ 1 Introduction
Problem (1) is ubiquitous in fields of machine learning and control and encompasses problems such as independent component analysis Nishimori (1999), low-rank matrix completion Vandereycken (2013), training of orthogonally normalized neural networks Bansal et al. (2018), the control of rigid bodies Duong et al. (2024), as well as sensor network localization Patwari & Hero (2004) and many others. Compared to general constrained optimization, Riemannian optimization promises the advantage of exploiting the natural geometry of the problem, producing feasible iterates and increased numerical robustness Boumal (2023).
@@
In contrast to the above classical setup, in this work we focus on the setting where the manifold $\mathcal{M}$ is given implicitly through a measure $\mu=\mu_{\mathrm{data}}$ supported on $\mathcal{M}$. This perspective is especially relevant in view of the manifold hypothesis (Loaiza-Ganem et al., 2024), which posits that many real-world data sets lie (approximately) on a manifold with dimension much smaller than that of the ambient space Fefferman et al. (2016). Importantly, such data manifolds are not just low-dimensional geometric structures—they also capture rich semantic meaning. For instance, the image manifold corresponds to photo-realistic images Pope et al. , the system behavior manifold to dynamically feasible input-output trajectories Willems & Polderman (1997), while the manifold of airfoils represents aerodynamically viable shapes Zheng et al. (2025). The optimization problem (1) in this setting thus encompasses a broad class of modern tasks, including airfoil design (Chen et al., 2025), reinforcement learning (Lee & Choi, 2025), and Bayesian inverse problems (Chung et al., 2022b). Driven by the growing demand in generative AI and data-driven design, this calls for a paradigm shift: moving from optimization over explicitly known manifolds, where classical Riemannian optimization applies, to optimization over data manifolds that are accessible only implicitly through samples. In this data-driven regime, methods from classical Riemannian optimization cannot be applied directly, since they rely on explicit manifold operations such as tangent-space projection, retraction, or exponential maps (Boumal et al., 2019; Boumal, 2023).
While there has been extensive research on manifold learning Meilă & Zhang (2024); Lin & Zha (2008); Cayton et al. (2005); Belkin & Niyogi (2005) in the past, none of these works addressed (1) from an optimization point of view and focused instead on learning the manifold geometry. Moreover, the non-parametric nature of these methods limits their applicability to very low-dimensional manifolds and cannot exploit the successful inductive bias and the exceptional representation power of modern neural networks.
@@
In this work, we propose a data-driven approach to recover the fundamental operations needed for optimization on manifolds. Starting from the data distribution $\mu_{\mathrm{data}}$, we smooth it with a Gaussian kernel to obtain
@@
$\ell_{\sigma}(x)=\frac{1}{2}\lVert x\rVert^{2}-\sigma^{2}\log p_{\sigma}(x).$ (3)
@@
We show that, as the smoothing parameter $\sigma$ decreases, the gradient $\nabla\ell_{\sigma}$ recovers the projection back to the manifold, while the Hessian $\nabla^{2}\ell_{\sigma}$ recovers the projection onto its tangent space. These results reveal that core ingredients of Riemannian optimization—such as retraction and gradient computation—can be implemented directly from the derivative information of $\ell_{\sigma}$, even when the manifold itself is only given implicitly through data.
@@
Given the above novel theoretical findings, a key practical challenge is how to access (even approximately) the gradient $\nabla\ell_{\sigma}$ and Hessian $\nabla^{2}\ell_{\sigma}$ when only samples from $\mu_{\mathrm{data}}$ are available. A crucial observation is that $p_{\sigma}$ coincides with the marginal distribution of the Variance-Exploding SDE (VE-SDE), with analogous arguments for VP-SDE and DDPM formulations (Ho et al., 2020; Song et al., 2020). In this setting, the gradient of $\log p_{\sigma}$ – commonly referred to as the score function – plays a central role. In diffusion models, this score is parameterized by a neural network and learned directly from samples of $\mu_{\mathrm{data}}$, typically via denoising score matching (Vincent, 2011; Song et al., 2020). It is therefore natural to adopt the same methodology here: a neural network can be trained to represent $\nabla\ell_{\sigma}$, while the Hessian $\nabla^{2}\ell_{\sigma}$ can be recovered by computing its Jacobian. This approach offers two key advantages: (i) strong inductive biases can be incorporated through the neural network parameterization, and (ii) efficient, well-established training techniques from the diffusion model literature can be directly leveraged. Taken together, the theoretical link between $\ell_{\sigma}$ and manifold geometry, and the practical machinery for learning the score, form the foundation of the paradigm shift: from classical Riemannian optimization with explicit manifold knowledge to a data-driven framework where geometry is recovered from samples—thus enabling principled manifold optimization in generative and design-driven applications.
Building on these insights, we propose two algorithms for optimization over data manifolds – denoising landing flow (DLF) and denoising Riemannian gradient descent (DRGD) – to the best of our knowledge, the first such framework in the literature. The prerequisite is access to an approximate function $s_{\theta}^{\sigma}\simeq\nabla\ell_{\sigma}$, which can be learned from data using standard diffusion model parameterization and training given samples from $\mu_{\mathrm{data}}$.
  • Our first method directly employs $s_{\theta}^{\sigma}$ and its Jacobian (with $\sigma$ fixed as a small constant) to substitute the corresponding manifold operations in Riemannian gradient descent: each iteration computes a Riemannian gradient step and then applies retraction to (approximately) return to the manifold.
  • Since learning the score is inevitably imperfect when $\sigma$ is small, we also design a complementary landing-type algorithm that avoids staying close to the feasible set at every step. Instead, it augments the target function $f$ with the regularization term $\ell_{\sigma}$ and performs gradient descent on this regularized objective.
@@
For both algorithms, we establish non-asymptotic convergence guarantees to approximate stationary points, measured by small Riemannian gradient norm, as $\sigma\to 0$. Importantly, our methods require only inference of the neural network and gradients with respect to its inputs—not with respect to the network parameters. Thus, if a pretrained score network is already available for a given task, no additional training is required to enable Riemannian optimization on the corresponding data manifold. Viewed from this perspective, our approach can also be interpreted as an inference-time algorithm, aligning with a growing trend in modern machine learning research.
@@ 1.1 Our contributions
We summarize our contributions and give the outline of the paper as follows.
  • Data-driven recovery of manifold operations. By introducing the link function $\ell_{\sigma}$, we develop strategies to recover core manifold operations directly from data. In Section 3, by leveraging the score function from diffusion models to train and parameterize $\nabla\ell_{\sigma}$ and $\nabla^{2}\ell_{\sigma}$, we bridge the gap between the requirements of classical Riemannian optimization (which assumes explicit manifold knowledge) and the emerging need to optimize over data manifolds that are only implicitly available.
  • First algorithms for optimization over data manifolds. We propose two algorithms – denoising landing flow (DLF) via (8) in Section 4 and denoising Riemannian gradient descent (DRGD) via (12) in Section 5to the best of our knowledge, the first in the literature – that exploit operations enabled by a pretrained score function. Our methods require only inexpensive inference-time queries and back-propagation with respect to the input of the neural network, making them computationally efficient and readily applicable when pretrained scores are available.
  • Rigorous non-asymptotic guarantees. We provide the first non-asymptotic convergence analysis for optimization over data manifolds. Our main results (Theorem 3 and Theorem 5) ensure both approximate feasibility (output is close to the manifold) and approximate optimality (small Riemannian gradient norm). A key technical contribution is a novel non-asymptotic analysis of how $\nabla\ell_{\sigma}$ and $\nabla^{2}\ell_{\sigma}$ approximate the true manifold operations when $\sigma$ is small but non-zero (Theorem 1).
Finally in Section 6 we validate our findings by numerically simulating the proposed flow on well-known manifolds from Riemannian optimization literature and data-driven optimal control for reference trajectory tracking. Our finding are that we can generate feasible points on the manifold with objective values far lower than the ones available in the training set – see Figure 1 – and shows the effectiveness of exploiting the strong inductive biases of modern deep learning for the classical field of constrained optimization.
@@ 1.2 Related Work
Riemannian Optimization.
Riemannian optimization (RO) was originally developed under the assumption that the constraint manifold is explicitly known, either through closed-form descriptions such as matrix manifolds or via nonlinear equality constraints (Sato, 2021; Boumal, 2023). A prototypical algorithm in this setting is Riemannian gradient descent, which updates by taking a gradient step along the tangent space and then retracting back to the manifold (Boumal et al., 2019). Such algorithms guarantee that the iterates remain on the manifold at every iteration and are thus referred to as feasible methods. In contrast, infeasible methods, where the iterates are not constrained to stay on $\mathcal{M}$, have also been studied—for example, augmented Lagrangian approaches (Xie & Wright, 2021). However, these typically involve solving complicated optimization subproblems in each inner loop, making them computationally expensive in practice. More recently, a new class of landing-type algorithms has emerged: instead of enforcing feasibility at every step, they regularize the objective function with the distance to the manifold and then perform gradient flow or descent on this regularized objective. Such methods have demonstrated strong empirical performance, offering a promising alternative to classical approaches (Ablin & Peyré, 2022; Schechtman et al., 2023).
@@
High-dimensional data in modern machine learning often exhibit an intrinsic lower-dimensional structure. Such structure is of central importance: it enables more efficient representation and compression of data, facilitates interpretability by revealing meaningful semantic organization, and provides a foundation for designing algorithms that exploit geometry rather than ambient dimensionality. The task of uncovering this structure is commonly referred to as manifold estimation or manifold learning, with a large body of work devoted to this goal. Classical approaches include Isomap (Tenenbaum et al., 2000), Laplacian Eigenmaps (Belkin & Niyogi, 2003), and Locally Linear Embedding (LLE) (Roweis & Saul, 2000), Diffusion Map (Coifman & Lafon, 2006), among many subsequent developments, e.g. (Zhou et al., 2020). We defer the discussion on diffusion-model-related manifold learning literature to the next paragraph.
@@
Diffusion models have achieved remarkable success in generative modeling, where the goal is to generate new samples consistent with an underlying data distribution $\mu_{\mathrm{data}}$. A central ingredient of these methods is the learning of the score function—the gradient of the log-density of the diffused distribution $p_{\sigma}$ Tang & Zhao (2025); Song et al. (2020). In practice, the score is parameterized by a neural network whose architectural design has been extensively studied, and its training is carried out using well-established techniques such as denoising score matching (Song et al., 2020). This combination of principled theory and mature practice has made diffusion models one of the most effective tools for data-driven generative modeling.
@@
Recent work on the statistical complexity of diffusion models under the manifold hypothesis provides indirect evidence that these models capture geometric information about the underlying data manifold (Oko et al., 2023; Tang & Yang, 2024). In particular, the sample complexity required to learn the data distribution $\mu_{\mathrm{data}}$ depends only on the intrinsic dimension of the manifold, rather than on the ambient dimension.
@@
Across the literatures Classifier(-Free) Guidance in Diffusion Models (Dhariwal & Nichol, 2021; Ho & Salimans, 2022) and the Plug-and-Play Framework in Bayesian Inverse Problems (Venkatakrishnan et al., 2013; Laumont et al., 2022; Pesme et al., 2025; Graikos et al., 2022; Chung et al., 2022a), a unifying perspective is the task of sampling from a posterior distribution of the form
@@
where $r$ denotes a cost function to be minimized—corresponding to our objective ff—and $\alpha>0$ is a temperature-like parameter (Domingo-Enrich et al., 2024). In Classifier(-Free) Guidance, $r$ encodes a classifier signal (e.g., specified by a prompt) and in Bayesian inverse problems, $r$ is the negative log-likelihood of observations. Meanwhile, $p_{\mathrm{pre}}$ serves as a prior distribution, typically derived from a large-scale pretrained generative model, which is expected to capture the semantic structure of the data manifold. Sampling from $p_{\mathrm{post}}$ thus aims to balance semantic plausibility with low cost.

### v2

# arXiv:2509.23357v2[cs.LG] 28 Feb 2026

@@ Abstract
Under the data manifold hypothesis, high-dimensional data concentrate near a low-dimensional manifold. We study Riemannian optimization when this manifold is only given implicitly through the data distribution, and standard geometric operations are unavailable. This formulation captures a broad class of data-driven design problems that are central to modern generative AI. Our key idea is a link function that ties the data distribution to the geometric quantities needed for optimization: its gradient and Hessian recover the projection onto the manifold and its tangent space in the small-noise regime. This construction is directly connected to the score function in diffusion models, allowing us to leverage well-studied parameterizations, efficient training procedures, and even pretrained score networks from the diffusion model literature to perform optimization. On top of this foundation, we develop two efficient inference-time algorithms for optimization over data manifolds: Denoising Landing Flow (DLF) and Denoising Riemannian Gradient Descent (DRGD). We provide theoretical guarantees for approximate feasibility (manifold adherence) and optimality (small Riemannian gradient norm). We demonstrate the effectiveness of our approach on finite-horizon reference tracking tasks in data-driven control, illustrating their potential for practical generative and design applications.
@@ 1 Introduction
Problem (1) is ubiquitous in fields of machine learning and control and encompasses problems such as independent component analysis Nishimori (1999), low-rank matrix completion Vandereycken (2013), training of orthogonally normalized neural networks Bansal et al. (2018), the control of rigid bodies Duong et al. (2024), as well as sensor network localization Patwari and Hero (2004) and many others. Compared to general constrained optimization, Riemannian optimization promises the advantage of exploiting the natural geometry of the problem, producing feasible iterates and increased numerical robustness Boumal (2023), thus allowing for early termination and making it more suitable for real-time implementation.
@@
In contrast to the above classical setup, in this work we focus on the setting where the manifold $\mathcal{M}$ is given implicitly through a finite set of samples from an underlying population data distribution $\mu_{\mathrm{data}}$ supported on $\mathcal{M}$. This perspective is especially relevant in view of the data manifold hypothesis (Loaiza-Ganem et al., 2024), which posits that many real-world data sets lie (approximately) on a manifold with dimension much smaller than that of the ambient space Fefferman et al. (2016). Importantly, such data manifolds are not just low-dimensional geometric structures – they also capture rich semantic meaning. For instance, the image manifold corresponds to photo-realistic images Pope et al. , the system behavior manifold to dynamically feasible input-output trajectories Willems and Polderman (1997), while the manifold of airfoils represents aerodynamically viable shapes Zheng et al. (2025). The optimization problem (1) in this implicit setting thus encompasses a broad class of modern tasks, including airfoil- and ship hull design (Chen et al., 2025; Bagazinski and Ahmed, 2023), additive manufacturing Peng et al. (2023), reinforcement learning (Lee and Choi, 2025), and Bayesian inverse problems (Chung et al., 2022b). In this data-driven regime, methods from classical smooth Riemannian optimization cannot be applied directly, since they rely on explicit manifold operations such as tangent-space projection, retraction, or exponential maps (Boumal et al., 2019; Boumal, 2023). Graph-based Riemannian optimization has recently been proposed in (Wang et al., 2025), but the setup is quite different, as to be elaborated in Section 1.2. Furthermore, while there has been extensive research on manifold learning Meilă and Zhang (2024); Lin and Zha (2008); Cayton and others (2005); Belkin and Niyogi (2005) in the past, none of these works addressed (1) from an optimization point of view and focused instead on learning the manifold geometry, rather than representation models that are suitable to be included as a constraint in an optimization problem.
@@
Instead, in this work, we propose a novel data-driven approach to recover the fundamental operations needed for optimization on manifolds: Starting first from the population data distribution $\mu_{\mathrm{data}}$, we smooth it with a Gaussian kernel to obtain
@@
$\ell_{\sigma}(x)=\frac{1}{2}\lVert x\rVert^{2}+\sigma^{2}\log p_{\sigma}(x)\,,\quad x\in\mathbb{R}^{d}\,.$ (3)
@@
We show that, as the smoothing parameter $\sigma$ decreases, the gradient $\nabla\ell_{\sigma}$ recovers the projection back to the manifold, while the Hessian $\nabla^{2}\ell_{\sigma}$ recovers the projection onto its tangent space. These results reveal that core ingredients of Riemannian optimization – such as retraction and gradient computation – can be implemented directly from the derivative information of $\ell_{\sigma}$, which itself can be constructed from $\mu_{\operatorname{data}}$.
@@
Given the above novel theoretical findings, a key practical challenge is how to access (even approximately) the gradient $\nabla\ell_{\sigma}$ and Hessian $\nabla^{2}\ell_{\sigma}$, when, in real-world applications, only samples from $\mu_{\mathrm{data}}$ are available. A crucial observation is that $p_{\sigma}$ coincides with the marginal distribution of the Variance-Exploding SDE (VE-SDE) [^1] Analogous extensions to VP-SDE and DDPM can also be similarly derived.. Note that the gradient of $\log p_{\sigma}$, i.e. the score function, is central to diffusion models. It is parameterized by a neural network and learned directly from samples of $\mu_{\mathrm{data}}$, typically via denoising score matching (Vincent, 2011). We can hence exploit this well-developed toolkit for a practical implementation of our idea: Let $\texttt{s}(x,\sigma)$ be a pretrained score network. We use $\texttt{v}(x)=x+\sigma^{2}\texttt{s}(x,\sigma)$ to approximately represent $\nabla\ell_{\sigma}$, while the Hessian $\nabla^{2}\ell_{\sigma}$ can be recovered by computing its Jacobian. This approach offers two key advantages: (i) strong inductive biases can be incorporated through the neural network parameterization, and (ii) efficient, well-established training techniques from the diffusion model literature can be directly leveraged. Taken together, the theoretical link between $\nabla\ell_{\sigma}$ and manifold geometry, and the practical machinery for learning the score, form the foundation of the paradigm shift: from classical smooth Riemannian optimization with explicit manifold knowledge to a data-driven framework where geometry is recovered from samples—thus enabling principled manifold optimization in generative and design-driven applications.
@@
Building on these insights, we propose the first score-based framework for optimization over data manifolds. We derive two algorithms: denoising landing flow (DLF) and denoising Riemannian gradient descent (DRGD). Both rely on an approximate score network $\texttt{s}(x,\sigma)\approx\nabla\log p_{\sigma}$ learned from samples of $\mu_{\mathrm{data}}$ using standard diffusion-model training. DRGD uses the learned gradient and its Jacobian (for a small, fixed $\sigma$) to mimic Riemannian gradient descent with retraction, while DLF performs gradient flow on a penalized objective, relaxing feasibility at intermediate iterates. We establish non-asymptotic convergence to approximate stationary points as $\sigma\to 0$. Importantly, our methods require only inference of the neural network and gradients with respect to its inputs—not with respect to the network parameters. Thus, if a pretrained score network is already available for a given task, no additional training is required to enable Riemannian optimization on the corresponding data manifold. Viewed from this perspective, our approach can also be interpreted as an inference-time algorithm, aligning with a growing trend in modern machine learning research.
@@ 1.1 Our contributions
We next summarize our main contributions and outline the structure of the paper.
  • Link function and data-driven manifold operations. Building on the smoothed data distribution $p_{\sigma}$ and the associated link function $\ell_{\sigma}$ introduced above, we show in Section 3 that its gradient and Hessian recover, in the small-$\sigma$ regime, the projection onto the manifold and its tangent space. Further, with the help of score learning in diffusion model training, we bridge the gap between the requirements of classical Riemannian optimization (which assumes explicit manifold knowledge) and the emerging need to optimize over data manifolds that are only implicitly available.
  • First score-based algorithms for optimization over data manifolds. We propose two algorithms – denoising landing flow (DLF) in Section 4 and denoising Riemannian gradient descent (DRGD) in Section 5 – to the best of our knowledge, the first in the literature that exploit operations enabled by a pretrained score function. Our methods require only inexpensive inference-time queries and back-propagation with respect to the input of the neural network, making them computationally efficient and readily applicable when pretrained scores are available.
  • Non-asymptotic guaranties. We establish non-asymptotic convergence guarantees for both algorithms, showing approximate feasibility (outputs close to the data manifold) and approximate optimality (small Riemannian gradient norm) as $\sigma\to 0$ (Theorems 3 and 5). A key technical ingredient is a uniform control of how $\nabla\ell_{\sigma}$ and $\nabla^{2}\ell_{\sigma}$ approximate the ideal manifold operations at finite $\sigma$ (Theorem 1).
Finally, in Section 6 we validate our approach on classical Riemannian optimization benchmarks and a data-driven optimal control problem for reference trajectory tracking. We demonstrate that our methods can generate feasible points on the manifold with objective values significantly lower than those observed in the training data (Figure 1), illustrating how strong inductive biases of modern deep networks can be harnessed for constrained optimization.
@@ 1.2 Related Work
In this section we briefly give an overview over the related work. More connections of our approach to previous work can be found in Appendix A.
Smooth Riemannian Optimization.
Riemannian optimization (RO) was originally developed under the assumption that the constraint manifold is explicitly known, either through closed-form descriptions such as matrix manifolds or via nonlinear equality constraints (Sato, 2021; Boumal, 2023). A prototypical algorithm in this setting is Riemannian gradient descent, which updates by taking a gradient step along the tangent space and then retracting back to the manifold (Boumal et al., 2019). Such algorithms guarantee that the iterates remain on the manifold at every iteration and are thus referred to as feasible methods. In contrast, infeasible methods, where the iterates are not constrained to stay on $\mathcal{M}$, have also been studied – for example, augmented Lagrangian approaches (Xie and Wright, 2021). However, these typically involve solving complicated optimization subproblems in each inner loop, making them computationally expensive in practice. More recently, a new class of landing-type algorithms has emerged: instead of enforcing feasibility at every step, they regularize the objective function with the distance to the manifold and then perform gradient flow or descent on this regularized objective. Such methods have demonstrated strong empirical performance, offering a promising alternative to classical approaches (Ablin and Peyré, 2022; Schechtman et al., 2023).
Optimization over graph
Recently, Wang et al. (2025) propose a graph-based strategy for manifold denoising, i.e. given observations sampled from the data manifold, for a new query point, find the closest point on the data manifold. The mixed-order optimization algorithm in this work consists of two types of updates: (1) By extending the concept of tangent space in classical smooth manifolds to the graph setting, first-order steps can be taken over the tangent bundle graph; (2) By enumerating the points connected by the “zero-order edge”, zero-th order steps can be taken to avoid local minima and achieve a strong result of convergence to global minima. We highlight that their setup is quite different from ours: (i) Their method is purely non-parametric and hence cannot exploit the recent development of deep learning; (ii) The proposed mix-order algorithm is more like a discrete search method. As an example, their strong global convergence result is built on the existence of zero-order edges, which “teleport” the iterate out of the local minima automatically. This is not the typical behavior of continuous optimization method, either deterministic or stochastic; (iii) Their method is purely designed for the denoising task while our focus is for minimizing general smooth objectives over data manifold. We highlight that due to the requirement of their analysis, for different objectives, the zero-order edges need to be reconstructed to avoid local minima.
@@
High-dimensional data in modern machine learning often exhibit an intrinsic lower-dimensional structure. Such structure is of central importance: it enables more efficient representation and compression of data, facilitates interpretability by revealing meaningful semantic organization, and provides a foundation for designing algorithms that exploit geometry rather than ambient dimensionality. The task of uncovering this structure is commonly referred to as manifold estimation or manifold learning, with a large body of work devoted to this goal.
  • Non-parametric approaches include Isomap (Tenenbaum et al., 2000), Laplacian Eigenmaps (Belkin and Niyogi, 2003), and Locally Linear Embedding (LLE) (Roweis and Saul, 2000), Diffusion Map (Coifman and Lafon, 2006), among many subsequent developments, e.g. (Zhou et al., 2020; Yao et al., 2023).
  • Parametric manifold learning techniques such as generative adversarial networks (GAN) and various types of autoencoders (AE) are used to estimate data manifolds $\mathcal{M}$ by learning a map $\psi:\mathcal{Z}\to\mathcal{M}$ that parametrizes $\mathcal{M}$ through some lower-dimensional “latent space” $\mathcal{Z}$ and which can be interpreted as a learned coordinate chart on the former Goodfellow et al. (2014); Berahmand et al. (2024). In the case of GAN this map corresponds to the generator, while for AE it corresponds to the decoder. For both GAN and AE, it has been noted that $\psi$ is not always a local chart in the differential-geometric sense Lee et al. (2022) and an additional line of work considered enforcing the latter property for the decoder through regularization or architectural constraints Sorrenson et al. (2023); Kumar et al. (2020). Manifold learning flows (M-flows) Brehmer and Cranmer (2020) address this problem by parametrizing $\psi$ in such a way that $\psi$ is injective and $\psi\circ\psi^{-1}$ can be efficiently computed and have been used for manifold learning and density estimation. Having obtained a chart $\psi$, one can locally trivialize the problem equation 1 to $\min_{\mathcal{Z}}f\circ\psi$, variants of which have also been known under the name of latent space optimization Tripp et al. (2020); Notin et al. (2021). We provide a detailed comparison of this paragraph to our work in Appendix A.
Comparison with Pull-back Gradient Flow.
When the map $\psi$ is learned, it induces a pull-back metric on the latent space $\mathcal{Z}$. Smooth Riemannian optimization algorithms on $\mathcal{M}$ can then be equivalently formulated in $\mathcal{Z}$; for instance, Riemannian gradient flow on $\mathcal{M}$ corresponds to a pull-back gradient flow on $\mathcal{Z}$. While we are not aware of prior work that explicitly exploits this idea to address optimization over data manifolds, we provide a comparison with this approach in Section A.2 and highlight several nontrivial challenges that should not be overlooked.
@@
Diffusion models have achieved remarkable success in generative modeling, where the goal is to generate new samples consistent with an underlying data distribution $\mu_{\mathrm{data}}$. A central ingredient of these methods is the learning of the score function – the gradient of the log-density of the diffused distribution $p_{\sigma}$ Tang and Zhao (2025); Song et al. (2020). In practice, the score is parameterized by a neural network whose architectural design has been extensively studied, and its training is carried out using well-established techniques such as denoising score matching (Song et al., 2020). This combination of principled theory and mature practice has made diffusion models one of the most effective tools for data-driven generative modeling.
@@
Recent work on the statistical complexity of diffusion models under the manifold hypothesis provides indirect evidence that these models capture geometric information about the underlying data manifold (Oko et al., 2023; Tang and Yang, 2024). In particular, the sample complexity required to learn the data distribution $\mu_{\mathrm{data}}$ depends only on the intrinsic dimension of the manifold, rather than on the ambient dimension.
@@
Across the literatures Classifier(-Free) Guidance in Diffusion Models (Dhariwal and Nichol, 2021; Ho and Salimans, 2022) and the Plug-and-Play Framework in Bayesian Inverse Problems (Venkatakrishnan et al., 2013; Laumont et al., 2022; Pesme et al., 2025; Graikos et al., 2022; Chung et al., 2022a), a unifying perspective is the task of sampling from a posterior distribution of the form
@@
where $r$ denotes a cost function to be minimized, corresponding to our objective $f$, and $\alpha>0$ is a temperature-like parameter (Domingo-Enrich et al., 2024). In Classifier(-Free) Guidance, $r$ encodes a classifier signal (e.g., specified by a prompt) and in Bayesian inverse problems, $r$ is the negative log-likelihood of observations. Meanwhile, $p_{\mathrm{pre}}$ serves as a prior distribution, typically derived from a large-scale pretrained generative model, which is expected to capture the semantic structure of the data manifold. Sampling from $p_{\mathrm{post}}$ thus aims to balance semantic plausibility with low cost.